I am working with systems of Linear Equations. Confused. can you show me step by step how to solve
-2x + y = 1 and 3x + y = 1

Answer by unlockmath(1688)

Hello,
The best way to do this is multiply the equation by -1 so to cancel out the y's
-2x(-1) + y(-1) = 1(-1) which gives us:
2x-y=-1 Now add this to the second equation:
3x + y = 1
And we get:
5x=0 Divide both sides by 5 and we get:
x=0
Now plug this into either original equation to find out "y"
-2(0) + y = 1 Do the calculations and we have:
y=1
There we go the solutions will work in either equation now.
RJ
Check out a new book I wrote at:
www.math-unlock.com